# Visualization of Electronic Health Record Sequences at Scale

### v1

@@ 3.4 Optimization of ESS Visualization Programs
2 chunk_size = 100,000 # tuned according to the processing speed
@@
6 for patients_chunk in split(patients, chunck_size):
7 tree = parallelized_aggregate(chunk, events)
@@
11def parallelized_aggregate(patients_chunk, events):
12 patients_per_thread = splitIn(patients, number_of_threads)
13 trees = [None] * number_of_threads # will receive the threads trees
14 parallel for thread in range(number_of_threads)
15 tree = aggregate(patients_per_thread[thread], events)
@@
17 for tree in trees[1:]: # merge all the trees in the first one
@@
Listing 2: Progressive Parallelized ESS Visualization. The type Node and function aggregate are the same as in LABEL:lst:simple.
@@
4␣␣newNode:Thenodeatthecurrentiteration.newNode.childrenshouldkeepthesameorderasoldNode.children
@@
Listing 3: The Hysteresis Sort algorithm in Python.
@@
The visualization of the ParcoursVis’s aggregated tree is an Icicle tree; it is laid out horizontally, node width encoding average (or median) node duration and node height encoding frequency, sorted to help compare siblings and show the frequency distribution per node. The stability of the aggregated visualization between iterations mainly depends on two node parameters: the width and the order of siblings. Changing the node width across updates can slightly shift subtrees, but changing sibling order can strongly change the overall layout, producing instability interpreted as uncertainty by users. We want to avoid this instability if it is only an artifact of the progressive algorithm.
@@
With our progressive aggregation algorithm, nodes with close frequencies could switch positions between iterations, e.g., two nodes containing almost 20% of the paths each could swap back and forth across iterations due to slight variations in the distribution of patients. We designed a Hysteresis Sort algorithm (LABEL:lst:hysteresis) to prevent these changes when the frequency difference is irrelevant visually. It is specifically designed for the progressive visualization of trees that rely on sorting the siblings by size.
@@
Our Hysteresis Sort sorts children perfectly at the first iteration. For each new iteration, when the size of two adjacent nodes is almost the same, we consider the two sizes as equivalent and maintain their previous order, as shown in LABEL:lst:hysteresis. Almost the same means the frequencies differ by a small number $\epsilon$; we call it the inertia. Our algorithm ensures that the nodes are almost sorted with an error bounded by the inertia, according to the following equation:
$\forall\{i,j\},i<j\implies\textit{freq}(\textit{child}_{j})-\textit{freq}(\textit{child}_{i})\leq\epsilon$ (1)
@@
We sort children in descending order, bottom to top, as we want new children created between two progressive iterations to be inserted at the end of the list of siblings (on top). This is because, statistically, rare nodes appear later than frequent nodes in the progression. Our algorithm is a variation of bubble sort. Since Equation 1 does not fulfill the triangular inequality (if $a-b\leq\epsilon$ and $b-c\leq\epsilon$, we cannot be sure that $a-c\leq\epsilon$), we are forced to verify the whole list again when children are moved, using the variable hasPermuted on line 27 of LABEL:lst:hysteresis. The worst-case complexity of the algorithm is $O(n^{2})$, but since the children are mostly sorted, its actual complexity is linear, and the number of children is usually small.
We choose a default inertia small enough to be unnoticeable visually, except at the first level, where it represents 20 pixels in height on a 1080p monitor. Below this value, two nodes are considered equivalent in size. Users can change inertia interactively if desired.
4 Evaluation
One goal of ParcoursVis is performance and scalability to explore large-scale pathways, visualizing billions of events aggregated while remaining interactive. Our evaluations report scalability, stability, completion time, and speed metrics. Usually, speed is completion time divided by the number of items considered, but with PVA, speed and completion time are decoupled as users can make early decisions if a partial result is good enough. We, therefore, report on the time taken for the results to become good enough to make an accurate decision regarding basic tasks. In addition to our git repository, we provide all data analyses on an osf.io repository.
4.1 Data
@@
ParcoursVis reads a dataset of low-level events, translates it into high-level ones, and builds its aggregated representation: the prefix tree. Our evaluation relies on our original synthetic dataset of $2M$ patients treated for non-cancerous prostate adenoma. It uses five low-level event types—Alphabloc, Phyto, Fivealpha, Surgery, and Death—and translates them into ten high-level event types grouped into four higher-level categories, shown in Table 2.
@@
Table 2: The ten event types used in our example use case.
To measure the scalability of ParcoursVis, we generated four synthetic datasets with $\textit{size}\in\{4M,6M,8M,10M\}$ patients using the strategy described in the supplementary material; their characteristics are listed below. The aggregated tree characteristics of the original 2M patients dataset are in the first line.
@@
4M 73,621,135 5,441,380 926 1,109
6M 110,526,755 8,164,831 948 1,134
8M 147,300,619 10,884,766 966 1,155
10M 184,161,763 13,607,620 972 1,161
@@
Figure 7: Total computation time per size compared with five chunk sizes using 6 threads. $\textit{chunkSize}=\infty$ means non-progressive.
@@
Figure 8: Total computation time per size compared by # of threads in progressive (dotted) and non-progressive (plain) settings.
@@
Before measuring the stability of ParcoursVis, we analyzed the time to aggregate the dataset with our progressive algorithm (i.e., chunk by chunk) compared to treating the whole dataset at once, not progressively. This competitive analysis is standard for online algorithms [6]. The reported completion times do not consider the rendering of the tree nor the possible network delay of a back-end/front-end architecture.
4.2 Aggregation
@@
We expect our algorithm (LABEL:lst:aggregation) to be linear in the number of events and threads. We report the computation times it takes to aggregate them both progressively and non-progressively. Instead of aggregating patients over a quantum of time, we aggregate with a fixed number of sequences for better control and reproducibility [14]. We vary $\textit{chunkSize}\in\{100K,150K,200K,250K\}$ sequences. As we expect our algorithm to be linear in the number of low-level events, we use that metric as the size of the databases in our evaluation. We ran our evaluations on an Intel(R) Core(TM) i7-8700K CPU @ 3.70GHz that has 6 cores and 12 logical threads. We disabled simultaneous multithreading (SMT) to enforce that each thread is associated with a physical core. We measured the time using a monotonic wall-clock.
4.2.1 Results
@@
It takes from 3.75s with one thread (2.67M patients per second) to 0.75s with six threads (13M patients per second) to process our biggest synthetic dataset of 10M patients in a non-progressive environment (Figure 7). The progressive environment yields similar results but with an overhead that depends on the chunkSize parameter. Figure 7 shows that the computation time decreases when the chunk size increases. Figure 8 also shows that the overhead increases with the number of threads. For the mono-thread configuration, the overhead is negligible. We explain this overhead by the number of tree merges the algorithm does (see LABEL:lst:aggregation line 18), which depends on the average and the number of threads as follows: $\textit{merges}=(\textit{nbThreads}-1)\times\textit{nbPatients}/\textit{chunkSize}$.
For $\textit{chunkSize}=100K$, $\textit{nbPatients}=10M$, and $\textit{nbThreads}=6$, we merge the trees 500 times, each composed of $\approx 1,100$ nodes.
For the best-case scenario of $\textit{chunkSize}=250K$ and $\textit{nbThreads}=6$, each progressive iteration takes about 19.819.8ms (median) to complete (12.612.6M patients per second).
@@
Figure 8 shows that, for all numbers of threads, the computation time of our algorithm remains linear in the number of low-level events for both progressive and non-progressive environments. Plotting the total computation time per size compared to the number of threads, we find that our algorithm’s running time is inversely proportional to the number of threads. Looking at the computational speed per size compared to mono-thread computations, we find that the speed increases linearly with the number of cores. A linear least-squares regression gives a slope $\approx 0.83$ and $R^{2}>0.99$. A one-way ANOVA on the linear regression predictions gives a $\textit{p-value}>0.99$; the slope coefficients for all the chunk sizes are the same.
@@
4.3 Stability
[…]
We considered two external factors affecting the ordering of nodes. First, we avoid what Monroe et al. [29] call confetti visualization by only showing nodes representing more than a certain MinSize number of patients. We evaluate ParcoursVis along the MinSize variable: we output a node and its children only if $\textit{size}(\textit{node})\geq\textit{MinSize}$ with $\textit{MinSize}\in\{0,25,50\}$.
@@ 4.3 Stability
The second factor impacting the stability of the tree is the convergence of the intermediate progressive results, which is sensitive to the number of sequences each iteration processes; larger chunks will lead to faster convergence. The number of intermediate steps depends on the CPU speed, machine load, and the time quantum, which are hard to control. We, therefore, vary the $\textit{chunkSize}\in\{100K,150K,200K,250K\}$ to control the number of patients processed per iteration to evaluate when nodes stabilize while ensuring reproducibility for the evaluation.
4.3.1 Metrics of Stability
For Hysteresis Sort and RSort, we report two metrics for stability. We also compute a binary metric Bad_Statistics, true if the frequency of a node at $\textit{iteration}=i$ differs from more than 5% from the final result. We use a frequency of 0 for not-yet-registered nodes. We report the lowest possible $\textit{iteration}=i$ for both metrics.
@@ 4.3.1 Metrics of Stability
Stability by Depth A depth is considered stable at iteration $i$ if, below that depth, there is no change of order in the trees generated from $\textit{iteration}=i$ to the final result.
@@
Stability Per Node A node is considered stable if its parent is stable, if this node and its siblings have the same rank in their parent’s list from $\textit{iteration}=i$ to the final result. The root is always considered stable. If a node is not created at a given iteration, we consider that it is placed at the end (last rank) of the children list.
@@
We emphasize that since we compare nodes from $\textit{iteration}=i$ to the final result, our report on stability is similar to convergence.
Figure 9: Stabilization speed of Hysteresis Sort and RSort by depth. Iteration ratio (0 to 1) indicates dataset processing progress.
@@
Figure 10: How fast the Hysteresis Sort and RSort stabilize nodes by depth.
@@
Figure 11: The stability of the tree by depth and MinSize values. Each category has 44 chunkSize values. We jitter the points on the x-axis to avoid point overlap for different MinSize values. The figure highlights clusters of categories: the higher the MinSize, the sooner the tree is stabilized. MinSize has a stronger effect on Hysteresis Sort than on RSort.
4.3.2 Results

### v2

@@ 3.4.1. Optimizing Data Preprocessing
2 chunk_size = 100,000 # tuned to processing speed
@@
6 for patients_chunk in split_by(patients, chunck_size):
7 tree = parallel_aggregate(chunk, events)
@@
11@dataclass class Node:
12 # No change from definition in Listing 1 line 8
13 def merge(self, other):
14 self.count += other.count
15 self.agedist.update(other.agedist)
16 self.durationdist += other.durationdist
17 for (event, node) in other.children.items():
18 if event in self.children:
19 self.children[event].merge(node)
20 else:
21 self.children[event] = node
22
23def parallel_aggregate(patients_chunk, events):
24 thread_patients = split(patients, threadcount)
25 trees = [None] * threadcount # tree per thread
26 parallel for thread in range(threadcount)
27 tree = aggregate(thread_patients[thread], events)
@@
29 for tree in trees[1:]: # merge all in the first one
@@
Listing 2 PVA Parallelized ESS aggregation Algorithm. The type Node and function aggregate remain the same.
@@
[…]
The first optimization relies on the memory layout, which organizes the low-level event sequences in consecutive memory positions so the aggregation algorithm can access them sequentially. Modern hardware strongly optimizes this access pattern by prefetching memory caches along with sequential access (Drepper, 2007). The aggregation process takes a table of low-level events indexed per patient and outputs a prefix tree meant to be rendered. In EventFlow, transforming low-level events into high-level ones is expensive due to the rich set of possible transformations supported, as shown in the Supplemental Material. In ParcoursVis, we write the function ll_to_hl (low level to high level) directly in C++ for two reasons: expressive power and optimization. As explained in Section 2.1.1, the aggregation functions are specific to treatments and use cases (e.g., non-cancerous prostate adenoma, used as an example here), so we hard-code the rules, e.g., the definition of what is an “interruption” synthetic event in a specific treatment, and make them follow state-of-the-art recommendations. Still, we allow the analysts to change some parameters within meaningful limits during the exploration (see Figure 6). ParcoursVis rules are more complex than what EventFlow can express, yet rather simple using a regular programming language like C++. Our aggregation rules implementation is 190 lines long. Adapting it to another scenario, described in Section 3.2, took a few days and had a similar length.
ȷ Second, for the progressive aspect, the main loop performs the tree aggregation along with the progressive calculations of the distributions. As shown in Listing 2 line 6, the back end partially aggregates the tree with a certain number of sequences (a chunk) and sends it to the front end (function render in Listing 2 line 8). When the function returns, meaning the user interface received the data, it continues the aggregation for the next chunk. To determine the chunk_size parameter for a given iteration (kept constant in the pseudo-code of Listing 2), we measure the average speed (patients per second) of the last six iterations and multiply it by the desired latency (typically 2 s). We start with the pessimistic chunk size of 100,000 patients and converge to a chunk size adapted to the network latency and performance of the back end.
We rely on the patients being shuffled for faster convergence of node frequency and average duration, and to overcome order bias in the data (Rusu et al., 2024). Keeping patients in chronological order would initially reveal pathways that followed older medical protocols and delay the visualization of recent ones, biasing the early overview snapshots.
We optimize the resulting tree to the rendering client with three transformations: (1) data structures stored in each node optimized for fast updates are transformed to be easily processed by the visualizations, e.g., the list of event durations is transformed into a histogram, (2) small nodes below a specified threshold of patients (10–50) are trimmed, and (3) node data not used by the tree layout or the overview screen are not included to minimize the transferred tree size. In our applications, the final tree has around 2 k nodes; the trimmed version can be transferred in a few milliseconds to the front end through a web socket.
As we keep the computed tree in memory, we distinguish between two kinds of data transfer: always transfered, the trimmed tree and basic attributes to be visualized as distributions, and on-demand node-level details which are only sent when the user explicitly requests them by clicking a node. This separation minimizes bandwidth usage and ensures low latency, even in environments with poor connectivity (e.g., hospitals with congested WiFi). With our strategy, web transfer latency only incurs when the user asks for details on demand, typically on a few nodes, and is low even when nodes contain rich aggregated data.
[…]
Progressively updating the tree and merging nodes is fast because we use data structures (not specific to PVA) that are efficient to update during sequence aggregation, and efficient to merge during node merging.
3.4.3. Progressive Rendering ȷ
Rendering is done on the front end in two passes: first, it preprocesses the tree according to user-specified parameters, such as the maximum tree depth and event filters, generating a rendering tree. Then, it lays out the Icicle tree in SVG format that eventually appears on the screen. The first step rewrites the aggregated tree into a new one when only simple transforms are involved, saving back-end heavy work. Typically, to filter out node types or apply a hierarchy transform (explained in Section 3.2.2). This allows for a smoother user experience since a tree is usually composed of a few thousand nodes that are fast to process compared to the dataset that can contain millions of events.
To prevent the progressive algorithm from flooding the rendering, the front end notifies the back end to continue the aggregation process once it has fully displayed an updated tree. When the user changes any of the parameters used for aggregation, the aggregation algorithm restarts from the beginning, iteratively updating the rendering, as shown by the arrows at the right of Figure 7.
3.4.4. Stabilizing Rendering ȷ
@@
4␣␣newNode:Thenodeatthecurrentiteration.
@@
Listing 3 The Hysteresis Sort algorithm in Python.
@@
The visualization of the ParcoursVis’s aggregated tree is an Icicle tree; it is laid out horizontally, node width encoding the average node duration and node height encoding the frequency, sorted to help compare the sibling frequency distribution per node. The stability of the aggregated visualization between iterations mainly depends on two node parameters: the width and the order of siblings. Changing the node width across updates can slightly shift subtrees, but changing sibling order can strongly change the overall layout, producing instability interpreted as uncertainty by users. We want to avoid this instability if it is only an artifact of the progressive algorithm.
@@
With our progressive aggregation algorithm, nodes with close frequencies could switch positions between iterations, e.g., two nodes containing almost 20% of the sequences each could swap back and forth across iterations due to slight variations in the distribution of patients (e.g., 19.9% vs. 20.1% and vice versa). We designed a Hysteresis Sort algorithm (Listing 3) to prevent these changes when the frequency difference is irrelevant visually. It is specifically designed for the progressive visualization of trees that rely on sorting the siblings by size.
@@
Our Hysteresis Sort sorts children perfectly at the first iteration. For each new iteration, when the size of two adjacent nodes is almost the same, we consider the two sizes as equivalent and maintain their previous order, as shown in Listing 3. Almost the same means the frequencies differ by a small number $\epsilon$; we call it the inertia. Our algorithm ensures that the nodes are almost sorted with an error bounded by the inertia, according to the following equation:
(1) $\forall\{i,j\},i<j\implies\textit{freq}(\textit{child}_{j})-\textit{freq}(\textit{child}_{i})\leq\epsilon$
@@
We sort children in descending order, bottom to top, as we want new children created between two progressive iterations to be inserted at the end of the list of siblings (on top). This is because, statistically, rare nodes appear later than frequent nodes in the progression. Our algorithm is a variation of bubble sort. Since Equation 1 does not fulfill the triangular inequality (if $a-b\leq\epsilon$ and $b-c\leq\epsilon$, we cannot be sure that $a-c\leq\epsilon$), we are forced to verify the whole list again when children are moved, using the variable hasPermuted (Listing 3 line 24). The algorithm has a worst-case complexity of $O(n^{2})$, but since the children are mostly sorted, its actual complexity is quasi-linear, and the number of children is usually small (ten in our prostate adenoma use case). We choose a default inertia small enough to be unnoticeable visually, except at the first level, where it represents 20 pixels in height on a 1080 p monitor. Below this value, two nodes are considered equivalent in size. Users can change inertia interactively if desired.
4. Evaluation
One of the primary goals of ParcoursVis is to ensure performance and scalability for the exploration of large-scale patient pathways, visualizing billions of aggregated events while maintaining interactivity. Rather than reinventing prior approaches, ParcoursVis builds upon EventFlow (Monroe et al., 2013), extending its principles to operate at a much larger scale. Our evaluations confirm that we achieve comparable levels of efficiency and usability, even as we scale to data sizes several orders of magnitude larger. We report metrics on scalability: latency, completion time, and stability. Accordingly, we report the time required for partial results to become sufficient for accurate decision-making in basic tasks. We provide all data analyses via an osf.io repository. We also report on qualitative users’ feedback.
4.1. Data
@@
ParcoursVis reads a dataset of low-level events, translates it into high-level ones, and builds its aggregated representation: the prefix tree. Our evaluation relies on our original synthetic dataset of $2M$ patients treated for non-cancerous prostate adenoma. It uses five low-level event types—Alphabloc, Phyto, Fivealpha, Surgery, and Death—and translates them into ten high-level event types grouped into four higher-level categories, shown in 4(a).
@@
(a) The ten event types used in our example use case.
@@
4 M 73,621,135 5,441,380 926 1,109
6 M 110,526,755 8,164,831 948 1,134
8 M 147,300,619 10,884,766 966 1,155
10 M 184,161,763 13,607,620 972 1,161
(b) Dataset sizes used in our evaluation.
Table 4. Event types and numbers used in our experiments.
To measure the scalability of ParcoursVis, we generated four synthetic datasets with $\textit{size}\in\{4\textrm{M},6\textrm{M},8\textrm{M},10\textrm{M}\}$ patients using the strategy described in the supplementary material; their characteristics are listed in 4(b). The aggregated tree characteristics of the original 2 M patients dataset are in the first line.
@@ 10 M 184,161,763 13,607,620 972 1,161
(a) Total computation time per size with five chunk sizes using 6 threads. $\textit{chunkSize}=\infty$ means non-progressive.
@@
(b) Total computation time per size compared by # of threads in progressive (dotted) and non-progressive (plain) settings.
Figure 8. Scalability comparison across chunk sizes (a) and thread counts (b).:
@@
Before measuring the stability of ParcoursVis, we analyzed the time to aggregate the dataset with our progressive algorithm (i.e., chunk by chunk) compared to treating the whole dataset at once, not progressively. This competitive analysis is standard for online algorithms (Borodin and El-Yaniv, 1998). The reported completion times do not consider the rendering of the tree nor the possible network delay of a back-end/front-end architecture.
4.2. Aggregation
@@
We expect our algorithm (Listing 2) to be linear in the number of events and threads. We report the computation times it takes to aggregate them both progressively and non-progressively. Instead of aggregating patients over a quantum of time, we aggregate with a fixed number of sequences for better control and reproducibility (Fekete and Freire, 2020). We vary $\textit{chunkSize}\in\{100\textrm{k},150\textrm{k},200\textrm{k},250\textrm{k}\}$ sequences. As we expect our algorithm to be linear in the number of low-level events, we use that metric as the size of the databases in our evaluation. We ran our evaluations on an Intel(R) Core(TM) i7-8700K CPU @ 3.70GHz that has 6 cores and 12 logical threads. We disabled simultaneous multithreading (SMT) to enforce that each thread is associated with a physical core. We measured the time using a monotonic wall-clock.
4.2.1. Results
@@
It takes from 3.75 s with one thread (2.67 M patients per second) to 0.75 s with six threads (13 M patients per second) to process our biggest synthetic dataset of 10 M patients in a non-progressive environment (8(a)). This result, which is below the 10 s threshold introduced in Section 2.2.1, should still be interpreted with caution. If the back-end machine is heavily loaded or if the network experiences slowdowns, the response time may become unstable and quickly exceed the 10 s threshold. Additionally, datasets that cover a large country or involve a more complex pathology, encompassing larger patient cohorts, may also lead to exceeding this limit. The progressive environment yields similar results but with a time overhead that depends on the chunkSize parameter. 8(a) shows that the computation time decreases when the chunk size increases. 8(b) also shows that the overhead increases with the number of threads. For the mono-thread configuration, the overhead is negligible. We explain this overhead by the number of tree merges the algorithm does (see Listing 2 line 30), which depends on the average and the number of threads as follows: $\textit{merges}=(\textit{nbThreads}-1)\times\textit{nbPatients}/\textit{chunkSize}$. For chunkSize=100 k, nbPatients=10 M, and nbThreads=6, we merge the trees with $\approx 1,100$ nodes 500 times. For the best-case scenario of chunkSize=250 k and nbThreads=6, each progressive iteration takes about 19.8 ms (median) to complete (12.6 M patients per second).
@@
8(b) shows that, for all numbers of threads, the computation time of our algorithm remains linear in the number of low-level events for both progressive and non-progressive environments. Plotting the total computation time per size compared to the number of threads, we find that our algorithm’s running time is inversely proportional to the number of threads. Looking at the computational speed per size compared to mono-thread computations, we find that the speed increases linearly with the number of cores. A linear least-squares regression gives a slope $\approx 0.83$ and $R^{2}>0.99$. A one-way ANOVA on the linear regression predictions gives a $\textit{p-value}>0.99$; the slope coefficients for all the chunk sizes are the same.
@@
4.3. Stability
We now focus on the stability and convergence of ParcoursVis over the progressive iterations. We compare our Hysteresis Sort, described in Section 3.4.4, with the regular sort (called RSort in the remaining) that EventFlow uses, i.e., sorting nodes based on frequency. We configured Hysteresis Sort with $\textit{inertia}=\frac{20}{1080}\approx 0.02$ by default, which corresponds to 20 pixels for the root level of our visualization on a standard 1080 p monitor, and below one pixel for the deeper tree levels.
@@ 4.3. Stability
We considered two external factors affecting the ordering of nodes. First, we avoid what Monroe et al. (Monroe et al., 2013) call confetti visualization by only showing nodes representing more than MinSize number of patients. We evaluate ParcoursVis along the MinSize variable: we output a node and its children only if $\textit{size}(\textit{node})\geq\textit{MinSize},\textit{MinSize}\in\{0,25,50\}$. The second factor impacting the stability of the tree is the convergence of the intermediate progressive results, which is sensitive to the number of sequences each iteration processes; larger chunks will lead to faster convergence. The number of intermediate steps depends on the CPU speed, machine load, and the time quantum, which are hard to control. We, therefore, vary the $\textit{chunkSize}\in\{100\textrm{k},150\textrm{k},200\textrm{k},250\textrm{k}\}$ to control the number of patients processed per iteration to evaluate when nodes stabilize while ensuring reproducibility for the evaluation.
4.3.1. Metrics of Stability
For Hysteresis Sort and RSort, we report two metrics for stability:
@@ 4.3.1. Metrics of Stability
Stability by Depth A tree depth is considered stable at iteration $i$ if, below that depth, there is no change of order in the nodes generated from $\textit{iteration}=i$ to the final result.
@@
Stability Per Node A node is considered stable if its parent is stable, if this node and its siblings have the same rank in their parent’s list from $\textit{iteration}=i$ to the final result. The root is always considered stable. If a node is not created at a given iteration, we consider that it is placed at the end (lowest frequency, last rank) of its sibling list.
@@
We report the lowest possible $\textit{iteration}=i$ for both metrics. We emphasize that since we compare nodes from $\textit{iteration}=i$ to the final result, our report on stability is similar to convergence.
(a) Trees’ stabilization speed for Hysteresis Sort and RSort by depth.
Iteration ratio $[0,1]$ indicates dataset processing progress.
@@
(b) Stabilization speed of the nodes using Hysteresis Sort and RSort by depth.
@@
(c) Stability of the tree by depth and MinSize. Each category shows 4 chunkSize values. Points are jittered on the x-axis to avoid overlap. Larger MinSize leads to earlier stabilization, especially in Hysteresis Sort.
Figure 9. Stability analysis of Hysteresis Sort and RSort: (a–b) stabilization results by depth and node count, and (c) effect of MinSize on overall tree stability.:
4.3.2. Results